OpenStudy (anonymous):

Solve for x: 2|x − 3| + 1 = 7 x = 0, x = 6 x = −1, x = 6 x = 0, x = 9 No solutions

OpenStudy (anonymous):

will fan & give medal

OpenStudy (bee_see):

replace x with the choices given to you and see both x's give you 7

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

2|x − 3| + 1 = 7 2|x − 3| + 1 - 1 = 7 - 1 2|x − 3| = 7 - 1 2|x − 3| = 6 2|x − 3| / 2 = 6 / 2 |x − 3| = 3 x − 3 = 3 and x − 3 = -3 Can you finish??